OpenStudy (anonymous):

Input the equation of the given line in standard form. The line with m = 2/3 and passing through (1, 1)

OpenStudy (anonymous):

y = mx +c 1 = (2/3)1 + c 1 - (2/3) = c c = 1/3 equation y = (2/3)x + (1/3)